OpenStudy (askme12345):

find all the degree values that fit between 0 < theta < 360 3cos 2 theta + 8 sin theta +5 = 0

OpenStudy (anonymous):

cos2t=1-2*sin^2 t so, 3*(1-2sin^2 t)+8sin t +5=0 3-6sin^2 t +8sint +5=0 simplify and solve it for sint like it's your x, after - find your t (t=theta)

OpenStudy (anonymous):

3cos 2 theta + 8 sin theta +5 = 0, 3 [1-2 sin^2 theta] + 8 sin theta + 5 =0, -2 sin^2 theta + 8 sin theta +8 =0. let x = sin theta, -2x^2 + 8x + 8 =0 now u have a quadratic problem. solve it then you'll get sin theta = .... to find theta, solve the above equation :)

OpenStudy (anonymous):

sorry it's minus 6 for the coefficient of x^2 ;)